# Price Experimentation and Interference

### v3

@@ 8.1 Specifications
    $R_{0}(f_{0},f_{1},\vec{p},q)$ and $R_{1}(f_{0},f_{1},\vec{p},q)$ are both differentiable over their domain.
@@
    $R_{1}(f,f,\vec{p},q)/q=R_{0}(f,f,\vec{p},q)/(1-q)=R(f,\vec{p}).$
Let $f_{1}^{n}$ be a sequence such that $\lim_{n\to\infty}f_{1}^{n}=f_{0}$. We then write the canonical estimator $\widehat{GTE}_{\Pi}$ as follows:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\lim_{n\to\infty}\frac{\Pi_{1}(f_{0},f_{1}^{n},\vec{p},q)/q-\Pi_{0}(f_{0},f_{1}^{n},\vec{p},q)/(1-q)}{f_{1}^{n}-f_{0}}.$
This is equivalent to:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\Pi_{1y}(f_{0},f_{0},\vec{p},q)/q-\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q),$ (30)
where $\Pi_{0y}$ and $\Pi_{1y}$ represent the derivative of the control and treatment platform profit functions with respect to the second argument respectively.
We define the bias as $Bias_{\Pi}(f_{0},\vec{p},q)=GTE_{\Pi}(f_{0},\vec{p},q)-\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)$. Using Assumption 8.1 we get that $\Pi(f_{0},\vec{p},q)=\Pi_{1}(f_{0},f_{0},\vec{p},q)/q$. We can then use an analogous argument to our derivation of (11) to write bias as:
$Bias_{\Pi}(f_{0},\vec{p},q)=\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+\Pi_{1x}(f_{0},f_{0},\vec{p},q)/q.$ (31)
Using out revenue functions and equation (28) we can rewrite the preceding expressions to be more useful. First we rewrite bias as:
$Bias_{\Pi}(f_{0},\vec{p},q)=f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+f_{0}R_{1x}(f_{0},f_{0},\vec{p},q)/q.$
Finally, we rewrite $\widehat{GTE}_{\Pi}$ from equation (30) as follows:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=R_{1}(f_{0},f_{0},\vec{p},q)/q+f_{0}R_{1y}(f_{0},f_{0},\vec{p},q)/q-f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q).$
As before, there are two necessary and sufficient conditions for the naive estimator to have the wrong sign in a fee experiment. Note that as was true in the pricing problem, bias is always positive when firms test a small price increase. This means the general conditions for a sign flip to occur are $GTE(f_{0},\vec{p})\geq 0$, meaning the true treatment effect of a small price increase is non-negative. And $Bias(f_{0},\vec{p},q)\geq\widehat{GTE}(f_{0},\vec{p},q)$, meaning there is enough bias to cause the estimated treatment effect to be non-positive.
These two conditions can be expressed as follows. First:
[…]
And second:
[…]
@@ 9 Conclusion and Future Work
In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue. We then show a novel method for debiasing canonical A/B experiments, by splitting half of subjects into treatment and half into control.
In future work we aim to extend our debiasing result to allow for correcting canonical experiments beyond the equal split between control and treatment. We also aim to run empirical studies to get a better sense of the prevalence of changes of sign in practice.

### v4

@@ 8.1 Model Calibration
    We set the total mass of listings $\rho=1$. This allows us to interpret $s^{*}(\lambda,p)$ as the fraction of listings that are available.
@@
    As usual, we normalize the value of the outside option, $\varepsilon=1$.
  3. 3.
    We set the cost $c=\hat{c}^{(s)}$.
  4. 4.
    We choose a reference price $p$ at which we want to evaluate price experiments. Actual prices were not reported in Farronato and Fradkin, (2022); in lieu of this information, in our results we vary $p$ in a “reasonable” range, as we describe below.
  5. 5.
    We choose an arrival rate $\lambda$. Again, no proxy for $\lambda$ is reported in Farronato and Fradkin, (2022), and so in our results we also vary $\lambda$ in a “reasonable” range, again as we describe below.
  6. 6.
    Finally, given the chosen values of $\rho,\ \varepsilon,\ p$ and $\lambda$, we choose $V$ so that the demand elasticity $pD^{\prime}(p)/D(p)$ for $D(p)=\rho-s^{*}(\lambda,p)$, with $s^{*}(\lambda,p)$ defined in (28), is equal to $\hat{e}_{p}^{(s)}$ (the estimated elasticity from Farronato and Fradkin, (2022)). It is straightforward to check the price elasticity of demand in our demand model is monotonically increasing in $V$, which guarantees that the fitted value of $V$ is uniquely determined in this approach.
Observe that at the conclusion of this procedure, for each value of $\lambda$ and $p$, we have an imputed demand model. We then use this demand model to determine whether a local CR or LR price experiment yields the wrong sign on the estimate of $GTE_{\pi}$, as discussed throughout the paper.
The two “free parameters” in our calibration are the reference price $p$, and the arrival rate $\lambda$, so we vary them in our subsequent empirical analysis. We vary the reference price $p$ in a plausible range from the cost $c$ to a value above the monopoly optimal price implied by $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. In our results we also vary $\lambda$ in a plausible range; in particular, we vary $\lambda$ so that in our fitted model we obtain plausible occupancy rates. The occupancy rate is the fraction of total supply that is booked in equilibrium, defined as $\zeta=1-s^{*}(\lambda,p)/\rho$. We vary $\lambda$ to ensure that we plot occupancy rates from 5% to 95%. As an example, one estimate from external market research on short term rental properties Airdna, (2024) estimates that the average occupancy rate across the entire Airbnb market is approximately 60%, with rates closer to 65% for major cities. This suggests the range of values of $\zeta$ in our results includes realistic values.
In our results, we present our visualizations with the two axes of price $p$ and occupancy rate $\zeta$, the latter in place of the arrival rate $\lambda$. We numerically verified that in this model, each pair of original parameters ($\lambda,p$), yields a unique occupancy pair ($\zeta,p$). Further, we have verified that in all instances of our model, as $\lambda$ increases, the resulting $\zeta$ increases monotonically as well. We make this transformation because the occupancy rate is a more natural metric for interpretation than the arrival rate of customers.
8.2 Results
Figure 5: Illustrations of the change of sign region (purple) in LR-experiments (top row) and CR-experiments (bottom row), respectively, for economy listings (left column) and luxury listings (right column) in Boston. The vertical axis is the imputed occupancy rate, and the horizontal axis is the (baseline) control price around which an experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue. We plot prices from the cost estimate $\hat{c}^{(s)}$, to a fixed amount above the critical price $p_{c}^{(s)}$.
Figure 5 shows the results of our model’s estimates for LR (top row) and CR (bottom row) price experiments in Boston. The left two plots show results for the economy tier, while the right two plots show results for the luxury tier. We have chosen to include Boston as an example city, but it is important to note that the trends observed in Figure 5, hold for all segments. (See Appendix 11 for these additional figures for all other segments.)
As in our preceding visualizations, the shaded blue region in each plot is where Condition (a) holds in the fitted demand model, cf. (13), and the shaded red region is where Condition (b) holds, cf. (14). These regions overlap in the purple region where both conditions hold. These are the regions where the estimated treatment effect would have the wrong sign. As noted in the preceding section, our results are plotted with the reference price $p$ on the x-axis (varied in a reasonable range, as described in the previous subsection), and the imputed occupancy rate $\zeta$ on the y-axis.
First, observe that in each figure, the right boundary of the change of sign region is a vertical line at a critical price $p_{c}^{(s)}$. To see why, observe that Condition (a) in (13) holds with equality at price $p_{c}^{(s)}$, with the cost $\hat{c}^{(s)}$ and elasticity $\hat{e}_{p}^{(s)}$ corresponding to the segment $s$. In particular, $p_{c}^{(s)}$ satisfies the first order condition for optimal monopoly pricing given $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. Further, this identity is insensitive to the value of the occupancy rate $\zeta$ (or equivalently, the arrival rate $\lambda$). Given the monotonicity in Condition (a), at any higher (resp., lower) values of $p$ than $p_{c}^{(s)}$, Condition (a) is guaranteed to not hold (resp., is guaranteed to hold); and again, these facts are also insensitive to the value of $\zeta$.
More importantly, observe that the figures make clear that changes of sign are prevalent in this model. Notably, across the parameter values we tested, LR experiments in particular seem especially vulnerable to changes of sign. This is consistent with the fact that the observed occupancy rates suggest a relatively demand-constrained market, at least as far as bias in price experiments is concerned. On the other hand, CR experiments exhibit a more nuanced behavior. For such experiments, a change of sign is observed in markets that are highly supply-constrained (e.g., with greater than 80-85% of listings being occupied), and when the reference price is near but still below the critical price $p_{c}^{(s)}$.
It should be noted that although CR price experiments exhibit more favorable behavior in these plots than LR price experiments, LR price experiments are more common in industry. In part this is because platforms are often reticent to offer individual customers different prices for the same product (Phillips, 2021). As a consequence, one potential practical takeaway from our work is that firms may in fact have chosen an experimental design that is more commonly producing the wrong sign in their estimated treatment effects in practice, at least in the settings we have studied. Despite the stylized nature of the model, the prevalence of changes of sign across a wide range of parameters suggests this is a phenomenon of significant consequence for practitioners.
@@ 9 Conclusion and Future Work
In this paper we studied the biases that arise due to interference in platform experiments, with a particular emphasis on profit and revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction. We then derive a novel method for debiasing canonical A/B experiments, by splitting units equally between treatment and control. We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
In future work we aim to extend our debiasing result, to allow for the correction of canonical experiments in cases where the treatment/control split is not 50/50. In addition to being more flexible, this may allow us to mitigate the sampling inefficiency of our current debiasing method. Currently, our method only uses half of the experimental data, namely, only the observed control demand, which may be a source of increased variance. Finally, we aim to run empirical studies to get a better sense of the true prevalence of changes of sign in practice.